# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=false, bottomline=false, skipabove=skipbelow=false

### v1

@@ Abstract
Data-driven software solutions have significantly been used in critical domains with significant socio-economic, legal, and ethical implications. The rapid adoptions of data-driven solutions, however, pose major threats to the trustworthiness of automated decision-support software. A diminished understanding of the solution by the developer and historical/current biases in the data sets are primary challenges.
To aid data-driven software developers and end-users, we present FairLay-ML, a debugging tool to test and explain the fairness implications of data-driven solutions. FairLay-ML visualizes the logic of datasets, trained models, and decisions for a given data point. In addition, it trains various models with varying fairness-accuracy trade-offs. Crucially, FairLay-ML incorporates counterfactual fairness testing that finds bugs beyond the development datasets. We conducted two studies through FairLay-ML that allowed us to measure false positives/negatives in prevalent counterfactual testing and understand the human perception of counterfactual test cases in a class survey. FairLay-ML and its benchmarks are publicly available at https://github.com/Pennswood/FairLay-ML. The live version of the tool is available at https://fairlayml-v2.streamlit.app/. We provide a video demo of the tool at https://youtu.be/wNI9UWkywVU?t=127.
@@ I Introduction
Addressing systemic bias and unfairness has been a fundamental societal problem, and nothing is more ”systemic” in the 21st century than life-altering decisions made by automated algorithms that often learn their logic from historical data. These automated decision-making algorithms act as gatekeepers for many social decisions: should a person remain in jail [1], is a person qualified for a job [2], and will a loan receiver become a default [3]. This is especially concerning when many of these algorithms are becoming less human-driven and human-understandable [4, 5]. The problem of human understandability and explainability particularly exacerbate discrimination in the data-driven automated algorithms since different stakeholders (e.g., software developers, domain experts, end users, etc.) increasingly find it more difficult to debug unfairness—the process of detecting, localizing, and fixing fairness bugs. This paper presents FairLay-ML, an intuitive debugging tool that aids data-driven software developers and end users to understand the fairness implications of their solutions.
@@
There are plenty of software tools in the literature to improve fairness [6, 7, 8, 9, 10]. IBM AIF360 [8] and Fairlearn [9] provide metrics and algorithms to measure unfairness and mitigate it via various pre-processing (e.g., modifying training datasets), in-processing (e.g., modifying the loss objective), and post-processing (e.g., calibration of scores) mechanisms. Google What-if toolkit [11] is the closest to FairLay-ML. In particular, both tools focus on individual fairness rather than group fairness, that is to ensure similar individuals receive similar outcomes. Google What-if toolkit allows to find a discriminatory instance by comparing a data point to its similar counterfactual who has a different protected attribute. FairLay-ML, on the other hand, goes beyond the training and development set and adapts individual fairness testing methods [12, 13, 14] to explore the arbitrary input space to find discriminatory instances. In addition, FairLay-ML allows users to modify test cases and define counterfactuals based on their intuitions (e.g., characteristics that make two individuals from different groups to be similar).
@@
We posit that significant user interactions in training and validating data-driven solutions against fairness requirements are critical design goals of the fairness toolkit. Hence, FairLay-ML integrates automation with human feedback, e.g., we adapt automated test-case generations, but also allow users to issue their own manual test cases based on their background and experiences. Throughout all the activities, we provide intuitive explanations and visualizations that guide users to make informed decisions.
@@
FairLay-ML is a GUI app that integrates multiple research ideas into the unified toolkit. The initial version is developed as a part of UG Thesis [15]. FairLay-ML takes a dataset of a target application (income, loan, credit, recidivism, etc.) as input and (i) visualizes the interactions between different features (e.g., distribution of job titles for male vs. female) via plotly and streamlit, (ii) runs a multi-objective evolutionary algorithm based on Parfait-ML [14] to train multiple models of varying accuracy vs. fairness based on a given threshold, (iii) explains the model logic based on the importance of different features on the outcome, (iv) conducts counterfactual study to generate test cases that may lead to individual discrimination, (v) explains how the model make a decision for a given test case, and (vi) allows users to create and modify a counterfactual for a given test case.
After providing the background, we describe the core design and implementation of FairLay-ML. Then, we overview the tool usage and showcase its usefulness for two empirical evaluations. In particular, we experiment to measure precision and recall of existing fairness testing techniques, thanks to the flexibility of FairLay-ML in finding and manipulating test cases. We also leverage FairLay-ML to conduct a class survey on the perception of counterfactuals by students. We discuss the limitations and conclude by describing future work.
@@ II Background
In the rapidly evolving field of Machine Learning (ML) [4, 5], fairness has become a main concern, specifically as ML software is being deployed or assisting in making high-stakes decisions such as hiring decisions [2], parole decisions [16], and financial lending [17]. These models often utilize large datasets that may contain historical or systematic biases to make decisions impacting people’s health, socio-economic, etc. Such biases in training data can lead the model to maintain or even amplify the existing disparities. This characteristic of ML systems may result in unfair outcomes based on legally protected attributes such as gender and race. For example, the recidivism risk assessment tool in criminal justice, COMPAS, has been under criticism for its biased predictions that disadvantage African Americans [16]. Amazon had to deactivate a hiring algorithm that excluded female applicants [2].
Fairness Notions. There are two broad fairness definitions: group fairness and individual fairness.
Group Fairness. It ensures that statistics of outcomes remain similar between two or more protected groups. The prevalent metric includes equal opportunity difference (EOD) which measures the differences between true positive rates between two groups and average odds difference (AOD) which measures the difference between two groups in the average of true positive and false positive rates [8, 18, 19]. Parfait-ML [14] used EDO and AOD to explore the hyperparameter space of ML algorithms (e.g., the depth of decision trees, the number of layers in DNNs, etc.) and find Pareto-optimal configurations that maximize fairness within a given acceptable accuracy range [14].
Individual Fairness. Fairness through awareness ($FTA$[20] requires two individuals who are deemed to be similar (based on their non-protected attributes) are treated similarly (regardless of their protected attributes). However, defining the notion of similarity between individuals is difficult. Counterfactual fairness [21] uses a causal dependency diagram between variables and requires that the ML outcome is causally independent of protected attributes. In this paper, we focus on a simpler notion, counterfactual discrimination as proposed by Galhotra et al. [12] that requires the ML outcome to remain similar between an individual and their counterfactual who has only a different protected value. This notion is the primary oracle of fairness testing [12, 22, 23, 13].
@@ III FairLay-ML
We present FairLay-ML, a GUI MLOps pipeline that allows users to upload their data and/or machine learning models to our toolkit for fairness debugging. FairLay-ML identifies, explains, and suggests remedies to improve the model fairness. We intend for FairLay-ML to be able to implement the remedies, retrain the model on the given dataset, and allow the user to re-evaluate the remedies for accuracy and fairness, as well as save the upgraded dataset and model, all within our pipeline. The key focus of FairLay-ML is counterfactual testing that measures the unfairness of data-driven software by computing the differences between an original test case and its counterfactual with different protected attributes. In doing so, FairLay-ML provides various data visualizations on how different attributes interact, training algorithms, and the logic of decision-making for a given data. The goal is to support users to investigate generated counterfactual test cases and manipulate counterfactuals based on their views.
III-A Feature Interactions
The data visualization tools provide an overview of how features correlate to the sensitive feature. This tool helps users explore the distribution of classifications for each feature, as well as understand how the features in a dataset are interconnected in the context of bias in machine learning. FairLay-ML ’s data visualization tool identifies key features that we should pay attention to in our analysis of bias. For example, zip code has been found in many contexts to accidentally be used by ML algorithms as a proxy for race [24], as can more obvious bugs in the pipeline such as “husband” and “wife” being used as a proxy for gender.
Our tool then seamlessly provides a user interface to dive deeper into the correlations for each feature. By clicking on the correlation graph directly, users are provided with a histogram of how likely a data point in each possible value of a specific feature is in a protected class of the sensitive feature. In doing so, users gain better context on whether specific values of a feature could be used when they modify an input to generate a counterfactual with a different protected attribute.
III-B Model Visualization and Explanation
Under the hood, FairLay-ML calls Parfait-ML to train different models of the same algorithm, varying accuracy vs. fairness. Then, FairLay-ML displays a model’s logic to gain (global) insights into how the model uses different features to make decisions. Currently, FairLay-ML supports displaying decision tree and random forest models by plotting each tree structure (or tree structure of all the trees in the random forest) directly, as well as logistic regression and support vector machine by graphing the weights of each feature or the slope of the hyperplane with respect to each feature respectively. Finally, FairLay-ML explains how a model classifies specific data points using local interpretable model-agnostic explanations called LIME [25]. LIME shows how each feature value of a data point impacts the model’s decision. This enables users to understand what features and what values may change the outcomes (e.g., from a favorable to an unfavorable one). In addition, these methods of visualization combine with feature interactions to provide explanations of the form “This model places a high weight on feature category X, and X is strongly correlated with the sensitive feature, causing unfairness.”
III-C Counterfactual Test Case Generation
Given a model and a set of test data points for the model, FairLay-ML can also scatterplot the data points for their predicted probability and highlight counterfactual points. This is useful during testing to see which individuals are specifically impacted by unfairness. For example, some models may only generate counterfactuals on the decision boundary (when the predicted probability is close to 0.5). Some models may have a bigger spread of counterfactuals, indicating a wider systemic gender bias within the model.
To dive deeper into a specific data point, users can filter by counterfactuals, and even if the counterfactual generated a more accurate or less accurate prediction. Users can also filter by false positive, false negative, true positive, or true negative. FairLay-ML then takes advantage of the previously mentioned explanation techniques such as LIME and model visualization to generate a common story for the user.
IV Tool Usage
@@
Figure 1 shows the end-to-end usage of FairLay-ML. A user begins by uploading their tabular data (both in the original form with labeled categories and features, as well as its numerically encoded) and a training algorithm (e.g., decision trees, logistic regressions, etc.). The user then specifies the sensitive attribute and categories within the sensitive attribute that the user wishes to study. From the data set, we first create an 80/20 train-test-split to have training and validation data sets. FairLay-ML also provides histograms and correlation metrics between each column and the sensitive attribute on the training data (feature interactions). FairLay-ML uses Parfait-ML [14] to generate various models with varying fairness/accuracy trade-offs. The visualization allows users to investigate different models where FairLay-ML plots the logic of models (e.g., normalized weights of contributions from different features). FairLay-ML also highlights Pareto-optimal models with red dots—those that achieve an ideal accuracy vs. fairness trade-off. Once the user picks one model, as the decision-making model, then FairLay-ML performs counterfactual testing based on the counterfactual discrimination notion using Themis [12]. The tool plots the generated test cases and highlights four types of data points: i) the original and counterfactual data points agree on a positive outcome; ii) the original and counterfactual data points agree on a negative outcome; iii) the original data point classified as positive, while the counterfactual one is classified as negative; and iv) the original data point classified as negative, while the counterfactual one classified as positive. The user can then investigate different data points. Once they click on one data point, FairLay-ML shows the local explanations on how the model outcomes a particular decision using LIME [25]. FairLay-ML also provides a dashboard to the users who can manipulate the data point to see how different changes contribute to the outcome. This enables users to investigate counterfactual test cases as well. They can use the LIME plot and identify which features/values might lead to a realistic counterfactual test case with significant discrimination.
V EMPIRICAL EVALUATION
@@
We performed a preliminary evaluation of FairLay-ML for debugging counterfactual test cases based on the following research questions (RQs):
@@
  • RQ2. How accurate are humans in generating counterfactual test cases?
Currently, FairLay-ML supports four datasets: Adult Census Income [26], Credit Bank [27], Loan Default [28], and COMPAS [29] over three protected attributes (sex, race, age). Since Adult Census Income is the most intuitive and simple applications (i.e., predicting whether incomes are above \$50k based on personal characteristics like the job, hours-per-week, sex, race, etc.), we will focus on this dataset to answer the research questions.
@@
RQ1. False Positives/Negatives of Counterfactual Testing. FairLay-ML implements a counterfactual testing method based on Themis [12]. The approach is to find individual discrimination (ID) instances—two inputs that share the same relevant inputs, but they receive two different outcomes due to their differences in the protected attributes—in a given amounts of time. Figure 2 shows 8,000 test cases visualized by FairLay-ML. The application is the Adult Census Income and sex is the protected attribute. The red points show ID instances where diamonds and square show whether the original instance or counterfactual one received a favorable outcome (an income above \$50k). However, simply perturbing protected attributes from $A$ to $B$ can lead to false positives or negatives. For example, an ID instance generated by changing the protected attribute from Male to Female may be false without adjusting the relationships (e.g., Husband to Wife). Similarly, a benign test case may be buggy once all the relevant features are adjusted for it. We study the false positive and negative rates of prevalent counterfactual testing. The key feature of FairLay-ML is that it allows to pick a test case and adjust any relevant features to generate a valid and proximal counterfactual. Thus, we use FairLay-ML to evaluate false positive and false negative rates of Themis approach. We pick 100 test cases uniformly at random from the set of test cases. The selected test cases are reported in [30]. We observe that the rate of false positives, false negatives, true positives, and true negatives are 21%, 16%, 37%, and 26%. An example of false positive is <sex="Male", race="Black", occupation="Handles-Cleaner", relationship="Husband", hours_per_week=80, native_country="Unknown"> where changing the sex attribute to ”Female” flips the outcomes from favorable (more than \$50k) to unfavorable (less than \$50k); but once the relationship is adjusted for (”Husband” $\to$ ”Wife”), then both original and counterfactual outcomes are favorable.
@@
RQ2. Validity and Proximity of Human Generated Counterfactuals. We perform a class survey with $17$ participants (UG and graduate students from the UT El Paso) using FairLay-ML. The details of survey can be found at [31] (permitted by the IRB). Figure 3 shows the steps to complete the survey where students can change up to 7 features of the target test cases (after changing the sex attribute from male to female) and report a combination that flips the ML decision from a favorable outcome (for the male individual) to an unfavorable one (for the female individual). In doing so, the students need to ensure that they preserve the validity and proximity of counterfactuals. The survey asked each student to evaluate the CFs reported by one or two other students in terms of its proximity to the target test case (with Id=25) and validity w.r.t real-world constraints. The results showed that 9 students out of 17 were able to find a combination that changes the ML outcome with a different sex and relationship. The number of changes vary between 2 and 7. The nine valid responses received 12 feedback on the validity and proximity. All the responses said ”yes” to the validity of CFs, implying that the counterfactual adheres to the real-world constraints. However, 6 of those evaluation feedback said that the CFs are not proximal to the original target test case (e.g., one counterfactual changed the education level). Most students tried to change race to black and believed that such changes may have significant impacts on the outcome, while being valid/proximal. We also observed that counterfactuals vary between participants significantly. For example, some participants said a separated individual may still have wife/husband relationship. Others believed that changing the country of origin does not preserve the proximity of CFs.
VI Limitations

### v2

@@ Abstract
Data-driven software solutions have significantly been used in critical domains with significant socio-economic, legal, and ethical implications. The rapid adoptions of data-driven solutions, however, pose major threats to the trustworthiness of automated decision-support software. A diminished understanding of the solution by the developer and historical/current biases in the data sets are primary challenges. To aid data-driven software developers and end-users, we present FairLay-ML, a debugging tool to test and explain the fairness implications of data-driven solutions. FairLay-ML visualizes the logic of datasets, trained models, and decisions for a given data point. In addition, it trains various models with varying fairness-accuracy trade-offs. Crucially, FairLay-ML incorporates counterfactual fairness testing that finds bugs beyond the development datasets. We conducted two studies through FairLay-ML that allowed us to measure false positives/negatives in prevalent counterfactual testing and understand the human perception of counterfactual test cases in a class survey. FairLay-ML and its benchmarks are publicly available at https://github.com/Pennswood/FairLay-ML. The live version of the tool is available at https://fairlayml-v2.streamlit.app/. We provide a video demo of the tool at https://youtu.be/wNI9UWkywVU?t=133.
@@ I Introduction
Addressing systemic bias and unfairness has been a fundamental societal problem, and nothing is more “systemic” in 21st century than life-altering decisions made by automated algorithms that learn their logic from historical data. These automated decision-making algorithms act as gatekeepers for many social decisions: should a person remain in prison [1], is a person qualified for a job [2], and will a loan receiver become a default [3]. This is concerning as many of these algorithms are becoming less human-driven and human-understandable [4]. The problem of human understandability and explainability exacerbates discrimination in data-driven automated algorithms since some stakeholders (e.g., software developers, domain experts, end users, etc.) can find it more difficult to debug unfairness—the process of detecting, localizing, and fixing fairness bugs. This paper presents FairLay-ML, an intuitive debugging tool that aids data-driven software developers and end users to understand fairness implications of their solutions.
@@
There are plenty of software tools in the literature to improve fairness [5, 6]. IBM AIF360 [5] and Fairlearn [6] provide metrics and algorithms to measure unfairness and mitigate it via pre-processing (e.g., modifying training datasets), in-processing (e.g., modifying the loss objective), and post-processing (e.g., calibration of scores) mechanisms. Google What-if toolkit [7] is the closest to FairLay-ML. Both tools focus on individual fairness rather than group fairness, that is to ensure similar individuals receive similar outcomes. Google What-if toolkit helps find a discriminatory instance by comparing a base data point to its similar counterfactual with a different protected attribute. FairLay-ML, instead, goes beyond the training and development set and adapts individual fairness testing methods [8, 9] to explore the arbitrary input space to find discriminatory instances. FairLay-ML also allows users to modify test cases and define counterfactuals based on their intuitions (e.g., characteristics that make two individuals from different groups similar).
@@
We posit that significant user interactions in training and validating data-driven solutions against fairness requirements are critical design goals. Hence, FairLay-ML integrates automation with human feedback, e.g., we adapt automated test-case generation, allowing users to issue manual test cases. Throughout all activities, we provide explanations and visualizations guiding users to make informed decisions.
@@
FairLay-ML is a GUI app that integrates multiple research ideas into a unified toolkit. FairLay-ML takes a dataset of a target application (income, loan, recidivism, etc.) as input and (i) visualizes interactions between different features (e.g., distribution of job titles by gender) via plotly and streamlit, (ii) runs a multi-objective evolutionary algorithm based on Parfait-ML [9] to train multiple models of varying accuracy vs. fairness based on a given threshold, (iii) explains the model logic based on the importance of different features on the outcome, (iv) generates base test cases that may lead to individual discrimination, (v) explains the model’s decision for a given test case, and (vi) allows users to create and modify counterfactuals of a given base test case.
@@ II Background
Fairness is an emergent concern as Machine Learning (ML) software is rapidly evolving, deployed to assist in high-stakes decisions such as in hiring [2], parole [10], and financial lending [11]. ML models are often trained with large datasets containing historical or systematic biases, and can lead models to maintain or amplify existing disparities, generating unfair outcomes based on legally protected attributes (e.g., gender, race). For example, the recidivism risk assessment tool, COMPAS, has yielded biased predictions that disadvantage African Americans [10] and Amazon had to deactivate a hiring algorithm that was biased against female applicants [2].
Notion of Individual Fairness. Fairness through awareness ($FTA$[12] requires two individuals deemed to be similar (based on their non-protected attributes) to be treated similarly (regardless of their protected attributes). However, defining the notion of similarity between individuals is difficult. Counterfactual fairness [13] uses a causal dependency diagram between variables and requires that the ML outcome is causally independent of protected attributes. Here, we focus on a simpler notion, counterfactual discrimination as proposed by Galhotra et al. [8] that requires the ML outcome to remain similar between an individual and their counterfactual who has only one different protected value. This notion is the primary oracle of fairness testing [8].
@@ III FairLay-ML
We present FairLay-ML, a GUI MLOps pipeline that allows users to upload data and/or machine learning models to our toolkit for fairness debugging. FairLay-ML identifies, explains, and suggests remedies to improve the model fairness. FairLay-ML implements the remedies, retrains the model on the given dataset, and allows the user to re-evaluate remedies for accuracy and fairness, and save the upgraded dataset and model, all within our pipeline. One key component of FairLay-ML is counterfactual testing that measures the unfairness of data-driven software by computing the differences between an original test case and its counterfactual with different protected attributes. In doing so, FairLay-ML provides various data visualizations on interactions between different attributes, relationships between configurations of training algorithms and fairness, and the logic of ML model in inferring a decision. This enables users to generate test cases (e.g., base and counterfactual test cases) that are natural and reflect their perceptions of real-world relationships.
@@
Figure 1 shows the end-to-end usage of FairLay-ML. A user begins by uploading their tabular data (both in the original form with labeled categories and features, as well as its numerically encoded) and a training algorithm (e.g., decision trees, logistic regressions, etc.). The user specifies the sensitive attribute and categories within the sensitive attribute they wish to study. From the dataset, we first create an 80/20 train-test-split to have training and validation datasets. FairLay-ML also provides histograms and correlation metrics between each column and the sensitive attribute on the training data (feature interactions). FairLay-ML uses Parfait-ML [9] to generate various models with varying fairness/accuracy trade-offs. The visualization allows users to investigate different models where FairLay-ML plots the logic of models (e.g., normalized weights of contributions from different features). FairLay-ML highlights Pareto-optimal models with red dots—those that achieve an ideal accuracy vs. fairness trade-off. Once the user picks one model as the decision-making model, FairLay-ML performs counterfactual testing based on the counterfactual discrimination notion using Themis [8]. The tool plots the generated test cases and highlights four types of data points: i) the original and counterfactual data points agree on a positive outcome; ii) the original and counterfactual data points agree on a negative outcome; iii) the original data point classified as positive, while the counterfactual one is classified as negative; and iv) the original data point classified as negative, while the counterfactual one classified as positive. The user can then investigate different data points. Once they click on one data point, FairLay-ML shows the local explanations on how the model generates a particular decision using LIME [14]. FairLay-ML presents a dashboard to users who can manipulate the data point to see how different changes contribute to the outcome, enabling users to investigate counterfactual test cases. They can use the LIME plot and identify which features/values might lead to a realistic counterfactual test case with significant discrimination. Next, we overview the key components of FairLay-ML.
Feature Interactions. The data visualization tools provide an overview of how features correlate to the sensitive feature, helping users explore the distribution of classifications for each feature, and understand how the features in a dataset are interconnected in the context of ML bias. FairLay-ML ’s data visualization tool identifies key features to pay attention to when analyzing bias. For example, zip code has often been found to accidentally be used by ML algorithms as a proxy for race, as can more obvious bugs in the pipeline such as “husband” and “wife” being used as a proxy for gender.
Our tool provides a user interface to dive deeper into correlations for each feature. By clicking on the correlation graph directly, users are shown a histogram of how likely each possible value of a specific feature to appear in a protected class of a sensitive feature. For example, Figure 1 includes ‘Feature Interactions’ where the attribute sex is “female”, the histogram of “husband” relationship has no density. This gives users better context on whether specific values of a feature could be used when they modify a test case.
Model Visualization and Explanation. FairLay-ML displays models’ logic to gain (global) insights into how features are used to make decisions. FairLay-ML supports displaying decision tree and random forest models by plotting all tree structure directly, while it shows the weights of each feature or the slope of the hyperplane with respect to each feature for other algorithms. Finally, FairLay-ML explains how a model classifies specific data points using local interpretable model-agnostic explanations called LIME [14]. LIME shows how each feature value of a data point impacts the model’s decision, enabling users to understand what features and values may change outcomes (e.g., from favorable unfavorable). These methods of visualization also combine with feature interactions to provide explanations of the form “This model places a high weight on feature category X, and X is strongly correlated with the sensitive feature, causing unfairness.” The ‘Select and Understand Model Logic’ box in Figure 1 shows how different features and their values contribute to the outcome of decision.
Counterfactual Test Case Generation. Given a model and a set of test data points for the model, FairLay-ML can scatterplot data points for their predicted probability and highlight counterfactual points to help during testing to see which specific individuals are impacted by unfairness. For example, some models may only generate counterfactuals on the decision boundary (when the predicted probability is close to 0.5). Some models may have a bigger spread of counterfactuals, indicating wider systemic bias.
Users can explore individual data points and filter specific counterfactuals even if it generated a more or less accurate prediction. Users can filter by false positive, false negative, true positive, or true negative. FairLay-ML then uses previously mentioned explanation techniques (e.g. LIME and model visualization) to generate a common story for the user. The ‘Manual Manipulation of Test Case’ box in Figure 1 shows an example of interactive counterfactual test case generation.
IV EMPIRICAL EVALUATION
@@
This preliminary evaluation of FairLay-ML for debugging counterfactual test cases asks these research questions (RQs):
@@
  • RQ2. How accurately do humans create counterfactuals?
Currently, FairLay-ML supports four datasets [15]: Adult Census Income, Credit Bank, Loan Default, and COMPAS over three protected attributes (sex, race, age). Since Adult Census Income is the most intuitive and simple applications (i.e., predicting whether incomes are above \$50k based on personal characteristics like the job, hours-per-week, sex, race, etc.), we will focus on this dataset to answer RQs.
@@
RQ1. False Positives/Negatives of Counterfactual Testing. FairLay-ML automatically generates test cases via a causal method, similar to Themis [8]. The approach is to find individual discrimination (ID) instances—two inputs that share the same relevant inputs, but they receive two different outcomes due to their differences in the protected attributes—in a given amounts of time. Figure 2 shows 8,000 test cases visualized by FairLay-ML. The application is the Adult Census Income and sex is the protected attribute. The red points show ID instances where diamonds and square show whether the original instance or counterfactual one received a favorable outcome (an income above \$50k). However, simply perturbing protected attributes from $A$ to $B$ can lead to false positives or negatives. For example, an ID instance generated by changing the protected attribute from Male to Female may be false without adjusting the relationships (e.g., Husband to Wife). Similarly, a benign test case may be buggy once all the relevant features are adjusted for it. We study the false positive and negative rates of prevalent counterfactual testing with only perturbing protected attributes. The key feature of FairLay-ML is that it allows to pick a test case and adjust any relevant features to generate a valid and proximal counterfactual. Thus, we use FairLay-ML to evaluate false positive and false negative rates. We pick 100 test cases uniformly at random from the set of test cases, and then manually evaluate their naturalness. We observe that the rate of false positives, false negatives, true positives, and true negatives are 20.4%, 27.5%, 72.5%, and 79.6%. An example of false positive is <sex="Male", race="Black", occupation="Cleaner", relationship="Husband", hours_week=80, country="Unknown"> where changing the sex attribute to “Female” flips the outcomes from favorable (more than \$50k) to unfavorable (less than \$50k); but once the relationship is adjusted for (“Husband” $\to$ ”Wife”), then both original and counterfactual outcomes become favorable.
@@
RQ2. Validity and Proximity of Human Generated Counterfactuals. FairLay-ML allows us to perform a study to understand the perception of counterfactuals by humans. We perform a class survey with $17$ participants (UG and graduate students from the UTEP) where they use FairLay-ML to modify a sample (e.g., the occupations from a private company to a non-profit organization) to generate a counterfactual with a different background based on their perception of validity (a natural counterfactual that adheres to real-world constraints) and proximity (a minimal number of changes to the original sample). Figure 3 shows the steps to complete the survey where students can change up to 7 features of the target test cases (after changing the sex attribute from male to female) and report a combination that flips the ML decision from a favorable outcome (for the male individual) to an unfavorable one (for the female individual) by querying the model using FairLay-ML. The survey asked each student to evaluate the CFs reported by one or two other students in terms of its proximity to the target test case and validity w.r.t real-world constraints. The results showed that 9 students out of 17 were able to find a combination that changes the ML outcome with a different sex and relationship. The number of changes vary between 2 and 7. The nine valid responses received 12 feedback on the validity and proximity. All the responses said ”yes” to the validity of CFs, implying that the counterfactual adheres to the real-world constraints. However, 6 of those evaluation feedback said that the CFs are not proximal to the original target test case (e.g., a counterfactual changed the education level). Most students tried to change race to black and believed that such changes may have significant impacts on the outcome, while being valid/proximal. We also observed that counterfactuals vary between participants significantly. For example, some participants said a separated individual may still have wife/husband relationship. Others believed that changing the country of origin does not preserve the proximity of CFs. FairLay-ML allowed i) students to easily make changes to a sample and query the model with arbitrary samples; and ii) authors to visualize the interactions between variables to evaluate validity/proximity of counterfactuals.
V Limitations